OpenStudy (anonymous):

Second order differential equations: 100(y'')-20(y)+y=0. Find general solution. Please explain how to obtain the general solution.

OpenStudy (anonymous):

PLEASE EXPLAIN STEPWISE!!

OpenStudy (anonymous):

isnt it 100 y'' -20 y' + y = 0

OpenStudy (anonymous):

oops yes

OpenStudy (anonymous):

ok you want a characteristic equation 100 r^2 - 20r + 1 = 0

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

you have a double root, r = 1/10

OpenStudy (anonymous):

ya...then?

OpenStudy (anonymous):

If tex2html_wrap_inline96 (which happens if tex2html_wrap_inline98 ), then the general solution is displaymath64

OpenStudy (anonymous):

the general solution is y = c1 *e^(1/10 x) + c2 *x *e^(1/10 x)

OpenStudy (anonymous):

is there a way to obtain without memorizing the formula?

OpenStudy (anonymous):

not that i know of

OpenStudy (anonymous):

there are 3 cases, distinct roots, double roots, and the imaginary roots. its not too bad, there is pattern

OpenStudy (anonymous):

the imaginary case is a little tricky

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Can this method also solve differential equations of the form y"=f(x)?

OpenStudy (anonymous):

how will u proceed after u take y''=f(x)?

OpenStudy (anonymous):

yes it usually makes it easier to reduce the quadratic, except in the case where there are radical linear factors

OpenStudy (anonymous):

show that the defferential equation y(y^2+2x)+2x(y^2+x)dy/dx=0 is not exact, but that it has an integrating factor of the form u=x^2y^k for some integer k. hence or otherwise find the general solution of this differential equation